# Integer Lattice Gas with a sampling collision operator for the fluctuating Navier-Stokes Equation

### v1

@@ II.2 Derivation of the Local Equilibrium Ensemble
This representation of the collision operator is entirely equivalent to the representation in Eq. (2). Due to the required momentum conservation the local equilibrium ensemble will, starting at $\pi=J$, increment by two, and thus will maintain its initial parity. Initial $N$- and $J$-values of opposing parity result in a maximum possible $\pi$-value of $N-1$. Also, since the particles occupying an $n_{0}$ state do not stream between lattice sites, and are only altered in multiples of two, the parity of the $n_{0}$-value at each lattice site from initialization acts as a spurious conserved quantity, the (minor) consequences of which will be discussed in the upcoming sections.
@@
where $k$ is an index which enumerates the possible values of $\pi$ while taking into account the fact that those values increment by multiples of two, and the floor function $\lfloor$..$\rfloor$ … ⌋ returns the greatest integer value less than its argument.
Figure 1: The Local equilbrium ensemble of $\pi$-values for $N=9$ and $J=0$. A single lattice site was made to undergo 100,000 collisions for both Blommel and Wagner’s algorithm and our sampling algorithm  Blommel and Wagner (2018). Then, a further 100,000,000 collisions were run, collecting the $\pi$-value after each. The histograms of these data sets are compared to the recursive and explicit formulae given by Eqs. (27) and (30) respectively
A comparison between this explicit formula and its recursive counterpart can be seen in Fig. 1, along with histograms of local equilibrium $\pi$-values from both Blommel’s algorithm and the sampling algorithm put forward in this paper. A low density system with a local $N$-value of $9$ was chosen due both to the fact that this local equilibrium ensemble is more distinct from a binomial distribution at lower densities and to the fact that we wished to display the lack of negative densities present in the distribution.
The histograms of data from both Blommel and Wagner’s algorithm and our sampling lattice gas algorithm show excellent agreement to the recursive and explicit formulae of the theoretical local equilibrium ensemble shown in Eqs. (27) and (30), as well as maintaining the proper parity for their initial $J$-value. To ensure proper agreement, additional tests were run for different configurations of $N$- and $J$-values, finding results consistent with those shown in Fig. 1. Thus, for brevity, we show only this single example.
@@
Blommel and Wagner followed Adhikari et al. Adhikari et al. (2005) in using arguments by Landau and Lifshitz (Landau and Lifshitz, 1969, Sec. 114) to postulate that the global equilibrium ensemble of a lattice gas should be given by a Poisson distribution around an equilibrium value Blommel and Wagner (2018). Due to the spurious parity conservation of the $n_{0}$-value at each lattice site this probability distribution cannot be the correct equilibrium ensemble for any specific set of initial conditions. Instead, particular microstates have to be removed from the total Poisson distribution. However, Blommel and Wagner showed numerically that the occupation numbers recovered in equilibrium are approximately Poisson distributed when the initial conditions themselves give a random distribution of even and odd $n_{0}$ and the equilibrium is understood as an ensemble average over these initial conditions.
The Poisson distribution of independent $n_{i}$-values around a given set of means $\tilde{f}_{i}$ is given by:
@@
where $P^{\mathrm{Pois}}(\{n_{i}\}_{l},\tilde{f}_{i})$ is the probability of finding a particular lattice site in state $\{n_{i}\}_{l}$. The $\tilde{f}_{i}$-values for Blommel and Wagner’s algorithm are given by Blommel and Wagner (2018),
@@
We can write Eq. (31) in terms of the system moments by utilizing Eqs. (9), (10), and (11). For brevity, the $n_{i}$-values are kept as $n_{i}(N,J,\pi)$ until otherwise required for calculation,
@@
We can further analyze this Poisson distribution by considering the requirements for equilibrium. Specifically, we require that our global ensemble must be invariant under Eq. (1). As the $n_{i}$-values from Eq. (31) are independently distributed, this distribution is inherently invariant under streaming. Therefore, it only remains to be proven that the system is invariant under collisions. We can test this analytically by considering the effect of a single collision on the probability distribution $P^{\mathrm{Pois}}[\{N,J,\pi\}_{l},f_{i}^{eq}(\bar{N}^{eq},\bar{J}^{eq})]$ as we did in Eq. (12), while adjusting for the fact that the effect of the collision operator is only required to be zero on average. Thus, we can write,
@@
We can further simplify this by taking the product over $i$, and substituting in Eqs. (9), (10), and (11) for the remaining $n_{i}$-values. However, for the sake of clarity, we first define factors $\alpha$ and $\beta$ as common terms on both sides of the equation after the substitution,
@@
Because Eq. (42) does not include all possible microstates of the Poisson distribution, it is no longer normalized. To account for this, Eq. (41) includes a normalization factor. This normalization factor consists of a sum over all possible values of $\pi$, utilizing an index $\pi^{\prime}$ to do so. However this representation does not take into account that the $\pi$-value can only be incremented by multiples of two. By defining a second index $k$, and setting $\pi^{\prime}=2k+|J|$, we can find a system that increases by two between values of $\pi^{\prime}$ with a lower bound of 00, and an upper bound of $\lfloor(N-|J|)/2\rfloor$. The floor function takes into account the fact that the maximum value of $\pi$ depends on the relative parity of $N$ and $J$.
Blommel and Wagner make the assumption when defining their collision rules that only local quantities can affect the local collisions Blommel and Wagner (2018). Thus, as a representation of the local equilibrium ensemble of Blommel and Wagner’s algorithm, Eq. (41) must only depend on local quantities. Thus, we must be able to rewrite Eq. (41) such that the dependency on the $f_{i}^{eq}$-values, which depend on the global quantities $\bar{N}^{eq}$ and $\bar{J}^{eq}$, vanishes. By using the aforementioned relation that $f_{1}^{eq}f_{-1}^{eq}/(f_{0}^{eq})^{2}=1/16$ to substitute $f_{0}=\sqrt{16f_{1}^{eq}f_{-1}^{eq}}$, we can rewrite $\gamma(\pi)$ as,
@@
A collection of these local equilibrium ensembles can then in theory form the proper global equilibrium ensemble for our system, taking into account the spuriously conserved $n_{0}$ parity by picking only $N$- and $J$-values for the global equilibrium ensemble that maintain this conservation.
@@ III The Boltzmann Limit and its Uniqueness
To derive the hydrodynamic limit of our lattice gas algorithm, which will be given by a set of continuity and Navier-Stokes equations, it is required that we first know the Boltzmann limit of our lattice gas. The Boltzmann limit of a lattice gas corresponds to the ensemble average over the possible microstates of a lattice gas system given a specific initial state and elapsed time. The Boltzmann limit of the lattice gas evolution equation then gives us the corresponding lattice Boltzmann equation. A Chapman-Enskog analysis (or just moment equation Wagner (1997)) of the lattice Boltzmann evolution equation Chapman and Cowling (1990) then recovers the continuity and Navier-Stokes equations of the corresponding lattice Boltzmann model.
For a lattice gas defined by the evolution equation given by Eq. (1), we can find its Boltzmann limit by taking the ensemble average over this evolution equation. We start by defining,
@@
where the non-equilibrium ensemble average delimited-⟨$\langle$..$\rangle$ … ⟩ is defined as,
@@
We can also derive the $f_{i}$-values for any system for which we know the ensemble average $N$-, $J$-, and $\pi$-values mentioned in the prior section,
@@
For our algorithm to have a unique Boltzmann limit, we must be able to derive an expression for $\Omega_{i}$ that depends only on the local $f_{i}$-values. For the Boolean lattice gas this is always possible if the $n_{i}$ are independently distributed, as the $f_{i}$-values then fully define the $n_{i}$ distribution Frisch et al. (1987); Broadwell and Han (2007). When velocity states are allowed to have more than a single particle occupying them, the $f_{i}$-values no longer fully define the $n_{i}$ distribution, even for uncorrelated particle distributions.
For example, we consider the probability distribution defined by a set of delta functions that on average lead to a specific set of $f_{i}$-values,
@@
$\prod_{i}(1-f_{i}+\lfloor f_{i}\rfloor)\delta(n_{i}-f_{i})+(f_{i}-\lfloor f_{i}\rfloor)\delta(n_{i}-f_{i}-1),$
which allows us to have a valid delta distributed $\{n_{i}\}_{l}$ state even for non-integer $f_{i}$-values. This ensemble is significantly different to the ensemble recovered from the Poisson distribution in Eq. (31), though their means remain identical.
A unique Boltzmann limit was found for the diffusive sampling lattice gas both by Chopard and Droz (Chopard and Droz, 1998, sec. 5.7) and by Seekins and Wagner Seekins and Wagner (2022). Both groups found that the diffusive sampling lattice gas recovered the BGK lattice Boltzmann collision operator in their Boltzmann limit, given by,
$\Omega_{i}^{\mathrm{BGK}}=-\omega\left(f_{i}-f_{i}^{0}\right).$ (56)
This lattice Boltzmann collision operator sets the particle distribution $\{f_{i}\}_{l}$ to some local equilibrium distribution $\{f_{i}^{0}\}_{l}$, at a rate given by the relaxation time, given by the inverse of our single-timestep particle collision fraction $\omega$. In the Boltzmann limit of the diffusive sampling lattice gas, $f_{i}^{0}$ is the polynomial lattice Boltzmann equilibrium distribution.
Blommel and Wagner showed that the BGK collision operator using the entropic local equilibrium distribution given by Eq. (32) is a good approximation in most cases Blommel and Wagner (2018). In order to test this for our sampling algorithm, we can compare the average collision operators for different distributions by looking at the difference in the average pre- and post-collision $\pi$-values that result from applying the collision operator to those initial distributions.
We also wished to consider the impact of correlations on the collision operator, as highly correlated systems may not have the same Boltzmann limit as uncorrelated systems. Specifically, we considered alternative sets of delta distributed particles with two highly correlated states. In both cases, either all particles at a given lattice site are moving, or all particles at that lattice site are at rest. The probability for a given lattice site to be in a moving state is given by the ratio of $\bar{\pi}$ to $\bar{N}$. For the first probability distribution, the moving particles are evenly distributed within a single lattice site, after taking the initial $\bar{J}$-value into account, giving us,
@@
As collisions require either a pair of rest particles or a pair of moving particles with opposing velocities, we can expect the number of collisions present in a system with a distribution given by $P_{\mathrm{corr2}}$ to be heavily diminished compared to other distributions, as well as only allowing for an increase in the total $\pi$-value of the system.
@@
To test the Boltzmann limit of our collision operator, we initialized four systems of 100,000 lattice sites utilizing the Poisson distribution given by Eq. (31), the set of delta functions given by Eq. (55), and the two highly correlated distributions $P_{\mathrm{corr1}}$ and $P_{\mathrm{corr2}}$. We started by defining constant (independent of $x$) initial $\bar{N}$- and $\bar{J}$-values and setting the initial $\bar{\pi}$-value to its minimum ($\bar{\pi}=0$). We were then able to define $f_{i}$-values corresponding to those average quantities using Eqs. (52), (53), and (54). Each system was then collided without streaming, collecting each pre- and post-collision set of $\pi$-values, and the average of the resulting difference between the two was taken to be our averaged collision operator.
@@
To compare with the lattice Boltzmann, the BGK collision operator was calculated directly utilizing the $\{f_{i}\}_{l}$ in Eq. (56) and with an equilibrium distribution given by Eq. (32). All simulations were then repeated, increasing the initial $\bar{\pi}$-value by 1 until the $\bar{\pi}$-value reached the $\bar{N}$-value selected.
@@
Figure 2: The averaged collision operator given an imposed average $\pi$-value for four lattice gas particle distributions compared to the BGK collision operator given an initial average ensemble of $\{\bar{N}=30,\bar{J}=0,\bar{\pi}=0\}$. Both the average collision operator (a) and the difference from the BGK calculated by subtracting the BGK from each collision operator (b) are shown. See text for further details.
@@
Fig. 2 shows the results of this test for a series of increasing average $\pi$-values. Fig. 2 (a) shows the average change in $\pi$-value over the course of a single collision, and Fig. 2 (b) shows the difference between each collision operator and the BGK collision operator.
@@
The highly correlated initial condition of Eq. (58) shows a behavior that is notably non-BGK. The equilibrium value of $\pi$ doesn’t correspond to a zero crossing for the collision operator, meaning that $\pi$alues already larger than the equilibrium value will initially grow for this distribution. Clearly this move away from equilibrium will change the distribution so that the $\pi$-value will eventually decrease towards the equilibrium value. However, it is quite unclear how such non-BGK behavior will affect the hydrodynamic transport parameters.
@@
Though the three uncorrelated collision operators and the collision operator resultant from $P_{\mathrm{corr1}}$ are similar, as can be seen in Fig. 2 (a), they are not identical, which Fig. 2 (b) shows more explicitly. We observe an oscillatory behavior of the delta distributed system. This is related to a slightly different average $\pi$-value for even or odd $n_{0}$. The slightly altered slope of the Poisson distributed system from the BGK and delta distributed systems can be interpreted as a slightly altered BGK relaxation value for $\omega$. Importantly here the equilibrium value for $\pi$ is unaffected.
@@
These results show that the Boltzmann limit of our algorithm is not unique. However, for the uncorrelated systems studied, the BGK entropic lattice Boltzmann algorithm is a good approximation. This entropic lattice Boltzmann algorithm is identical to the algorithm first developed by Ansumali et al. Ansumali et al. (2003).
@@
Using the previously discussed Chapman-Enskog analysis, the approximate continuity and Navier-Stokes equations for the D1Q3 sampling lattice gas can be derived as Ansumali et al. (2003),
@@
$\nu_{\omega}=\theta\left(\frac{1}{\omega}-\frac{1}{2}\right),$ (61)
@@
Blommel and Wagner show in Fig. 5 of their paper that this value for $\theta$ is only a good approximation for $|u|<0.4$ Blommel and Wagner (2018). Outside of that range, $\theta$ will have some dependence on the density, and thus the Navier-Stokes equation will have pressure and viscosity terms that are density dependent.
The kinematic viscosity in Eq. (61) can be lowered by selecting an $\omega$-value between $1$ and $2$, which is useful for simulating low-viscosity systems. This is easily done for the BGK collision operator of Eq. (56). These systems are generally referred to as over-relaxed. However there is no obvious way for the lattice gas to access over-relaxed collisions due to the fact that there is no obvious way to collide more than $100\%$ of particles at a given lattice site. Some work has been done on this problem by Strand and Wagner Strand and Wagner (2022), who created a method to simulate over-relaxation in diffusive systems, but we leave the extension of these methods to hydrodynamic systems to future work.

### v2

@@ II.2 Derivation of the Local Equilibrium Ensemble
This representation of the collision operator is entirely equivalent to the representation in Eq. (2). To put this in the context of continuous fluctuating methods: the alteration of the $\pi$ value in the collision operator represents a microscopic implementation of the fluctuating stress tensor. Due to the required momentum conservation the local equilibrium ensemble will, starting at $\pi=J$, increment by two, and thus will maintain its initial parity. Initial $N$ and $J$ values of opposing parity result in a maximum possible $\pi$ value of N1N-1. Also, since the particles occupying an $n_{0}$ state do not stream between lattice sites, and are only altered in multiples of two, the parity of the $n_{0}$ value at each lattice site from initialization acts as a spurious conserved quantity, the (minor) consequences of which will be discussed in the upcoming sections.
@@
where $k$ is an index which enumerates the possible values of $\pi$ while taking into account the fact that those values increment by multiples of two, and the floor function $\lfloor$$\rfloor$ returns the greatest integer value less than its argument.
Figure 1: The Local equilbrium ensemble of $\pi$ values for $N=9$ and $J=0$. A single lattice site was made to undergo 100,000 collisions for both Blommel and Wagner’s algorithm and our sampling algorithm  [19]. Then, a further 100,000,000 collisions were run, collecting the $\pi$ value after each. The histograms of these data sets are compared to the recursive and explicit formulae given by Eqs. (26) and (29) respectively
A comparison between this explicit formula and its recursive counterpart can be seen in Fig. 1, along with histograms of local equilibrium $\pi$ values from both Blommel’s algorithm and the sampling algorithm put forward in this paper. A low density system with a local $N$ value of $9$ was chosen due both to the fact that this local equilibrium ensemble is more distinct from a binomial distribution at lower densities and to the fact that we wished to display the lack of negative densities present in the distribution.
The histograms of data from both Blommel and Wagner’s algorithm and our sampling lattice gas algorithm show excellent agreement to the recursive and explicit formulae of the theoretical local equilibrium ensemble shown in Eqs. (26) and (29), as well as maintaining the proper parity for their initial $J$ value. To ensure proper agreement, additional tests were run for different configurations of $N$ and $J$ values, finding results consistent with those shown in Fig. 1. Thus, for brevity, we show only this single example.
@@
Blommel and Wagner followed Adhikari et al. [11] in using arguments by Landau and Lifshitz [28, Sec. 114] to postulate that the global equilibrium ensemble of a lattice gas should be given by a Poisson distribution around an equilibrium value [19]. Due to the spurious parity conservation of the $n_{0}$ value at each lattice site this probability distribution cannot be the correct equilibrium ensemble for any specific set of initial conditions. Instead, particular microstates have to be removed from the total Poisson distribution. However, Blommel and Wagner showed numerically that the occupation numbers recovered in equilibrium are approximately Poisson distributed when the initial conditions themselves give a random distribution of even and odd $n_{0}$ and the equilibrium is understood as an ensemble average over these initial conditions.
The Poisson distribution of independent $n_{i}$ values around a given set of means $\tilde{f}_{i}$ is given by:
@@
where $P^{\mathrm{Pois}}(\{n_{i}\}_{l},\tilde{f}_{i})$ is the probability of finding a particular lattice site in state $\{n_{i}\}_{l}$. The $\tilde{f}_{i}$ values for Blommel and Wagner’s algorithm are given by [19],
@@
We can write Eq. (30) in terms of the system moments by utilizing Eqs. (8), (9), and (10). For brevity, the $n_{i}$ values are kept as $n_{i}(N,J,\pi)$ until otherwise required for calculation,
@@
We can further analyze this Poisson distribution by considering the requirements for equilibrium. Specifically, we require that our global ensemble must be invariant under Eq. (1). As the $n_{i}$ values from Eq. (30) are independently distributed, this distribution is inherently invariant under streaming. Therefore, it only remains to be proven that the system is invariant under collisions. We can test this analytically by considering the effect of a single collision on the probability distribution $P^{\mathrm{Pois}}[\{N,J,\pi\}_{l},f_{i}^{eq}(\bar{N}^{eq},\bar{J}^{eq})]$ as we did in Eq. (24), while adjusting for the fact that the effect of the collision operator is only required to be zero on average. Thus, we can write,
@@
We can further simplify this by taking the product over $i$, and substituting in Eqs. (8), (9), and (10) for the remaining $n_{i}$ values. However, for the sake of clarity, we first define factors $\alpha$ and $\beta$ as common terms on both sides of the equation after the substitution,
@@
Because Eq. (41) does not include all possible microstates of the Poisson distribution, it is no longer normalized. To account for this, Eq. (40) includes a normalization factor. This normalization factor consists of a sum over all possible values of $\pi$, utilizing an index $\pi^{\prime}$ to do so. However this representation does not take into account that the $\pi$ value can only be incremented by multiples of two. By defining a second index $k$, and setting $\pi^{\prime}=2k+|J|$, we can find a system that increases by two between values of $\pi^{\prime}$ with a lower bound of 0, and an upper bound of $\lfloor(N-|J|)/2\rfloor$. The floor function takes into account the fact that the maximum value of $\pi$ depends on the relative parity of $N$ and $J$.
Blommel and Wagner make the assumption when defining their collision rules that only local quantities can affect the local collisions [19]. Thus, as a representation of the local equilibrium ensemble of Blommel and Wagner’s algorithm, Eq. (40) must only depend on local quantities. Thus, we must be able to rewrite Eq. (40) such that the dependency on the $f_{i}^{eq}$ values, which depend on the global quantities $\bar{N}^{eq}$ and $\bar{J}^{eq}$, vanishes. By using the aforementioned relation that $f_{1}^{eq}f_{-1}^{eq}/(f_{0}^{eq})^{2}=1/16$ to substitute $f_{0}=\sqrt{16f_{1}^{eq}f_{-1}^{eq}}$, we can rewrite $\gamma(\pi)$ as,
@@
A collection of these local equilibrium ensembles can then in theory form the proper global equilibrium ensemble for our system, taking into account the spuriously conserved $n_{0}$ parity by picking only $N$ and $J$ values for the global equilibrium ensemble that maintain this conservation.
@@ III The Boltzmann Limit and its Uniqueness
To derive the hydrodynamic limit of our lattice gas algorithm, which will be given by a set of continuity and Navier-Stokes equations, it is required that we first know the Boltzmann limit of our lattice gas. The task of obtaining a Boltzmann limit for a lattice gas goes back to the very origin of lattice gases [1]. Even in the continuous case of obtaining the Boltzmann equation from the BBGKY hierarchy [3] it is necessary to make the significant assumption that correlations can be neglected, and that the two particle distribution function can be factorized as the product of two one-particle distribution functions. Usually it is assumed that it is the pre-collision two-particle distribution function that is uncorrelated, and this approximation introduces the irreversibility absent in the full BBGKY hierarchy. It is understood that this approximation will only be approximately valid, and the same difficulty exists for the Boltzmann limit of our lattice gas, as we will see below.
To define the Boltzmann limit of a lattice gas, we consider the ensemble average over the possible microstates of a lattice gas system given a specific initial state and elapsed time. The above mentioned difficulty is then moved to the task of defining the appropriate ensemble over which to average. For the Boolean lattice gases the expectation value $\langle n_{i}\rangle$ already defines the distribution of the $n_{i}$ completely and it only remains to make the assumption $\langle n_{i}n_{j}\rangle=\langle n_{i}\rangle\langle n_{j}\rangle$ [1]. In the case of an integer lattice gas just knowing $\langle n_{i}\rangle$ does not imply the underlying distribution function, giving more freedom in the definition of the ensemble over which to average.
The Boltzmann limit of the lattice gas evolution equation then gives us the corresponding lattice Boltzmann equation. A Chapman-Enskog analysis (or just moment equation [29]) of the lattice Boltzmann evolution equation [30] then recovers the continuity and Navier-Stokes equations of the corresponding lattice Boltzmann model.
For a lattice gas defined by the evolution equation given by Eq. (1), we can formally find its Boltzmann limit by taking the ensemble average over this evolution equation. We start by defining,
@@
where the non-equilibrium ensemble average $\langle$$\rangle$ is defined as,
@@
We can also derive the $f_{i}$ values for any system for which we know the ensemble average $N$, $J$, and $\pi$ values mentioned in the prior section,
@@
For our algorithm to have a unique Boltzmann limit, we must be able to derive an expression for $\Omega_{i}$ that depends only on the local $f_{i}$ values. The difficulty that arises is that the $f_{i}$ do not uniquely define the $P(\{n_{j}\}_{g},t)$. The molecular chaos assumption usually applied can be written as
$P(\{n_{j}\}_{g},t)\approx\prod_{x}\prod_{i}P(n_{i},x,t)$ (54)
where the $P(n_{i},x,t)$ is the probability for the occupation number $n_{i}(x,t)$ to occur.
Now for the Boolean lattice gas knowing $f_{i}$ is sufficient to know the one particle distribution function, which is simply given by $P(n_{i}\equiv 1,x,t)=f_{i}$ and $P(n_{i}\equiv 0,x,t)=1-f_{i}(x,t)$ [1, 31]. When velocity states are allowed to have more than a single particle occupying them, the $f_{i}$ values no longer fully define the $n_{i}$ distribution, even for uncorrelated particle distributions, and we will consider the effect of this freedom below.
Given the general discussion above, it may seem surprising that a unique Boltzmann limit was found for the diffusive sampling lattice gas both by Chopard and Droz [16, sec. 5.7] and by Seekins and Wagner [21]. This case is subtly simpler, because collisions can be written as one-particle collisions with an imagined substrate, making correlations irrelevant in the derivation of the collision operator. Since all particles equally contribute to collisions, independent of their current velocity state, the form of the one-particle distribution is likewise irrelevant. Both groups found that the diffusive sampling lattice gas recovered the BGK lattice Boltzmann collision operator in their Boltzmann limit, given by,
$\Omega_{i}^{\mathrm{BGK}}=-\omega_{LB}\left(f_{i}-f_{i}^{0}\right).$ (55)
This lattice Boltzmann collision operator sets the particle distribution $\{f_{i}\}_{l}$ to some local equilibrium distribution $\{f_{i}^{0}\}_{l}$, at a rate given by the relaxation time, given by the inverse of $\omega_{LB}$, which is analogous to the particle collision fraction $\omega$ in a lattice gas. In the Boltzmann limit of the diffusive sampling lattice gas, $f_{i}^{0}$ is the polynomial lattice Boltzmann equilibrium distribution.
The full analytical derivation of the collision operator in the hydrodynamic case is an interesting problem. While we looked at this problem, and some progress appears to be possible, it is outside the scope of the current paper. Instead we will numerically evaluate the effect of different one particle distribution functions corresponding to the same set of expectation values $f_{i}$ on the collision operator, and we will see that this choice can indeed affect the collision operator. We also show that more significant effects are possible if the factorization assumption of Eq. (54) is dropped. This is important to make sense of some of our simulations shown below that resulted in a significant deviation from the BGK collision operator.
Blommel and Wagner showed that the BGK collision operator using the entropic local equilibrium distribution given by Eq. (31) is a good approximation in most cases [19]. To test this for our sampling algorithm, we can compare the average collision operators for different ensembles by looking at the difference in the average pre- and post-collision $\pi$ values that result from applying the collision operator to those initial distributions.
The first ensemble we consider is actually a fixed value for the $n_{i}$, given by $n_{i}=f_{i}$, when the $f_{i}$ happen to be integers and if they aren’t integers then we interpolate them such that the mean values recover the non-integer $f_{i}$ by defining
@@
$\prod_{i}(f_{i}-\lfloor f_{i}\rfloor)\delta(n_{i}-\lfloor f_{i}\rfloor+1)+(1-(f_{i}-\lfloor f_{i}\rfloor))\delta(n_{i}-\lfloor f_{i}\rfloor).$ (57)
This is what we refer to as the “Delta distributed $n_{i}$” below. The second distribution we consider is a Poisson distributed set of $n_{i}$ which defines the probability of such states as
$P^{Poisson}(\{n_{j}\})=\prod_{i}e^{-f_{i}}\frac{n_{i}^{f_{i}}}{n_{i}!}.$ (58)
This particular distribution would recover an equilibrium distribution when $\langle\pi\rangle$ corresponds to the equilibrium value. Each of these distributions fulfills our requirement that $\langle n_{i}\rangle=f_{i}$. We will see below that collisions applied to these ensembles are close to, but not identical to, the BGK collision operator.
In some of the simulations presented below we observed deviations that we were too large to be explained by different uncorreleated distributions, like those of Eqn. (57) and Eqn. (58). We therefore also wanted to consider correlated distributions. Specifically, we considered alternative sets of delta distributed particles with two highly correlated states. In both cases, either all particles at a given lattice site are moving, or all particles at that lattice site are at rest. The probability for a given lattice site to be in a moving state is given by the ratio of $\bar{\pi}$ to $\bar{N}$. For the first probability distribution, the moving particles are evenly distributed within a single lattice site, after taking the initial $\bar{J}$ value into account, giving us,
@@
As collisions require either a pair of rest particles or a pair of moving particles with opposing velocities, we can expect the number of collisions present in a system with a distribution given by $P_{\mathrm{corr2}}$ to be heavily diminished compared to other distributions, as well as only allowing for an increase in the total $\pi$ value of the system.
@@
To test the Boltzmann limit of our collision operator, we initialized four systems of 100,000 lattice sites utilizing the Poisson distribution given by Eq. (30), the set of delta functions given by Eq. (56), and the two highly correlated distributions $P_{\mathrm{corr1}}$ and $P_{\mathrm{corr2}}$. We started by defining constant (independent of $x$) initial $\bar{N}$ and $\bar{J}$ values and setting the initial $\bar{\pi}$ value to its minimum ($\bar{\pi}=0$). We were then able to define $f_{i}$ values corresponding to those average quantities using Eqs. (51), (52), and (53). Each system was then collided without streaming, collecting each pre- and post-collision set of $\pi$ values, and the average of the resulting difference between the two was taken to be our averaged collision operator.
@@
To compare with the lattice Boltzmann, the BGK collision operator was calculated directly utilizing the $\{f_{i}\}_{l}$ in Eq. (55) and with an equilibrium distribution given by Eq. (31). All simulations were then repeated, increasing the initial $\bar{\pi}$ value by 1 until the $\bar{\pi}$ value reached the $\bar{N}$ value selected.
@@
(b) . Both the average collision operator (a) and the difference from the BGK calculated by subtracting the BGK from each collision operator (b) are shown. See text for further details.
Figure 2: The averaged collision operator given an imposed average $\pi$ value for four lattice gas particle distributions compared to the BGK collision operator given an initial average ensemble of $\{\bar{N}=30,\bar{J}=0,\bar{\pi}=0\}$ with $\omega=\omega_{LB}=1$
@@
Fig. 2 shows the results of this test for a series of increasing average $\pi$ values. Fig. 2 (a) shows the average change in $\pi$ value over the course of a single collision, and Fig. 2 (b) shows the difference between each collision operator and the BGK collision operator.
@@
The highly correlated initial condition of Eq. (60) shows a behavior that is notably non-BGK. The equilibrium value of $\pi$ doesn’t correspond to a zero crossing for the collision operator, meaning that $\pi$ -alues already larger than the equilibrium value will initially grow for this distribution. Clearly this move away from equilibrium will change the distribution so that the $\pi$ value will eventually decrease towards the equilibrium value. However, it is quite unclear how such non-BGK behavior will affect the hydrodynamic transport parameters.
@@
Though the three uncorrelated collision operators and the collision operator resultant from $P_{\mathrm{corr1}}$ are similar, as can be seen in Fig. 2 (a), they are not identical, which Fig. 2 (b) shows more explicitly. We observe an oscillatory behavior of the delta distributed system. This is related to a slightly different average $\pi$ value for even or odd $n_{0}$. The slightly altered slope of the Poisson distributed system from the BGK and delta distributed systems can be interpreted as a slightly altered BGK relaxation value for $\omega_{LB}$. Importantly here the equilibrium value for $\pi$ is unaffected.
@@
These results show that the Boltzmann limit of our algorithm is not unique. However, for the uncorrelated systems studied, the BGK entropic lattice Boltzmann algorithm is a good approximation. This entropic lattice Boltzmann algorithm is identical to the algorithm first developed by Ansumali et al. [20].
@@
Using the previously discussed Chapman-Enskog analysis, the approximate continuity and Navier-Stokes equations for the D1Q3 sampling lattice gas can be derived as [20],
@@
$\nu_{\omega}=\theta\left(\frac{1}{\omega_{LB}}-\frac{1}{2}\right),$ (63)
@@
Blommel and Wagner show in Fig. 5 of their paper that this value for $\theta$ is only a good approximation for $|u|<0.4$ [19]. Outside of that range, $\theta$ will have some dependence on the density, and thus the Navier-Stokes equation will have pressure and viscosity terms that are density dependent.
The kinematic viscosity in Eq. (63) can be lowered by selecting an $\omega_{LB}$ value between $1$ and $2$, which is useful for simulating low-viscosity systems. This is easily done for the BGK collision operator of Eq. (55). These systems are generally referred to as overrelaxed. However for the lattice gas the collision probability $\omega$ has to be between zero and one. Some work has been done on this problem by Strand and Wagner [32], who created a method to simulate overrelaxation in diffusive systems, but their approach cannot be extended to hydrodynamic systems. Instead an additional step in the lattice gas approach has to be implemented, and even then the results cannot access the zero viscosity limit of $\omega=2$ for a finite number of particles per lattice cell. This extension, however, is outside the scope of this paper and will be addressed in a forthcoming publication.